OpenStudy (kl0723):

can someone help me understand why this is equal to e^2...Limit as n-> infinity of [1+(2/n)]^n

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

what is your definition of \(e^x\)?

OpenStudy (misty1212):

one is \[\lim_{n\to \infty}\left(1+\frac{x}{n}\right)^n\]

OpenStudy (misty1212):

in which case replacing \(x\) by \(2\) gives \(e^2\) right of the bat

OpenStudy (misty1212):

elsewise you have to take the log and use l'hopitals rule for the resulting limit you will get 2 so final answer will be \(e^2\)

OpenStudy (irishboy123):

start with the original identity (from Bernoulli compounding formula, the origin of e).... \[\lim_{n \rightarrow \infty} ( 1 + \frac{1}{n})^n = e\] then play with some substitution, eg let k = 2*n or n/2 and do the limit for k -> infinity

OpenStudy (kl0723):

Thank you!